OpenStudy (anonymous):

Which of the following functions increases, going upwards from left to right on a graph? A. y = 2(1/3)^x B. y = -2(3)^x C. y = 2x + 6 D. y = 3(1/2)^x

OpenStudy (anonymous):

@Luis_Rivera

OpenStudy (anonymous):

Which of the following functions decreases, going downwards from left to right on a graph? A. y = 4x B. y = 5x C. y = (1/2)^x D. y = 3(2)^x

OpenStudy (anonymous):

y did u delete ur answer

OpenStudy (anonymous):

I need confirm my answer since you give out the second one. I don't want you confused. let me check and give you the right one. sorry for that.

OpenStudy (anonymous):

the second one is C , for sure

OpenStudy (anonymous):

for the first question, I don't know why there is no solution satisfy the condition you give out. If B without minus sign. It comes true. Unfortunately, it's not. so I don't know.